OpenStudy (anonymous):

Convert the rectangular equation to polar form. Assume a>0. x^2+y^2 = 4a^2

OpenStudy (anonymous):

use r^2=x^2 + y^2 tan(theta) = y/x notice there's no more x's and y's to replace after replacing that left side with r^2.

OpenStudy (anonymous):

Thank you, I guess the part I don't really understand is what to do with the "a" on the right side. So with r^2 = 4a^2.

OpenStudy (anonymous):

well your equation can be r = 2a. this is a circle with radius 2a in polar form...

OpenStudy (anonymous):

Alright, so if I had an equation such as y=2a, I could take the y=r sin(). So it would be r sin() =2a. But is that in polar form, or do I have to do something else?

OpenStudy (experimentx):

just change \(x = r \cos \theta \) and \( y = r \sin \theta \)

OpenStudy (experimentx):

and simplify it!!